OpenStudy (anonymous):

is 98 a perfect square ? explain why

OpenStudy (anonymous):

Do you still need an answer?

OpenStudy (anonymous):

nope lol

OpenStudy (anonymous):

Okay. Lol

OpenStudy (anonymous):

I need another answer now

OpenStudy (anonymous):

lol but it might take a while

OpenStudy (anonymous):

Whats the problem?

OpenStudy (anonymous):

Explain how you could use the Pythagorean Theorem to find the distance between the points (-4,-3) and (1,1) . Find the distance. You can type "^2" to indicate "squared" and "sqrt" to indicate "square root." (Examples: 5^2=5^2 and sqrt (25)=√25

OpenStudy (anonymous):

Wow.. I'm going to be straight up with you, I have no clue

OpenStudy (anonymous):

lol I thought so lol

OpenStudy (anonymous):

it's fine ill someone else lol thanks

OpenStudy (anonymous):

Okay. Lol.